# Theory of enhanced-by-coincidence neural information transmission

### v1

@@ G.1 Sketch of the derivation of the temporal distribution function in the quenched-noise approximation
where $t,t^{\prime}:=\ell j,\ell j^{\prime}$, $P(\cdot)={\cal N}(\cdot;{\mu_{\rm I}},{v_{\rm c}})$ is the stationary distribution of input noises, and $P(\mu,t|\mu^{\prime},t^{\prime})$ is the Green’s function of an Ornstein-Uhlenbeck process: (mind $\xi$ in Eq. (5)):
@@
Both nested integrations, that in ${r_{\rm w}}(\mu_{t})\right>_{P(\mu_{t}|\mu_{t^{\prime}})}$ (compacting the notation) and, afterwards, that in
@@
can be performed in the saddle-point approximation, for large ${\tau_{\rm c}}/{\tau_{\rm m}}$. It is straightforward to arrive to an expression for ${g_{\rm q}}$ in the quenched-noise approximation. Interestingly, in the low- and large- $|t-t^{\prime}|$ limits, the ${r_{\rm R}}{g_{\rm q}}(t-t^{\prime})$ take, respectively, the values ${r_{\rm w}}^{2}(\mu)\right>_{P(\mu)}$ and ${r_{\rm w}}(\mu)\right>^{2}_{P(\mu)}$ (as expected, since for large $|t-t^{\prime}|$ we expect $\mu,\mu^{\prime}$ to be uncorrelated, and so ${r_{\rm w}}(\mu)$ and ${r_{\rm w}}(\mu^{\prime})$). In particular, for large enough $|t-t^{\prime}|$ (or, more precisely, neglecting terms of order $\omega^{2}$ in front of first-order terms in $\omega$), we get a simplified expression for ${g_{\rm q}}$:
@@
where $Q$ is a constant in time, depending on the white-LIF properties ${\mu_{\rm I}},{\sigma_{\rm I}},{\tau_{\rm m}}$, and on ${v_{\rm c}}$ (see the explicit expression in Eq. (61) below in Sec. G.2).
G.2 Evaluation of the nested integrals
@@
where ${t_{\rm w}}^{\prime},{t_{\rm w}}^{\prime\prime}$ and ${\sf g}$ are in Eq. (41,47), and where:
@@
Appendix H Hand-waving estimation for the rate variance in the quenched-noise approximation
H.1 Hand-waving ansatz for the temporal distribution function
We here present a hand-waving argument, consistent with the quenched-noise approximation, that leads to an alternative analytic expression for the rate variance of the LIF model with white+colored noise input. Interestingly, the hand-waving approximation for ${v_{\rm R}}$ and $g$ captures relatively well the behaviour of the numeric estimation of these quantities, but worst than the full quenched-noise calculation in Sec. G, of which the hand-waving approximation is a further simplification (please, see Figs. 6,7).
@@ H.1 Hand-waving ansatz for the temporal distribution function
${g_{\rm q}}^{(\text{hand})}(t)=\frac{{v_{\rm q}}}{{r_{\rm q}}}e^{-|t|/{\tau_{\rm c}}}+{r_{\rm q}},$ (69)
where the coefficients $a={r_{\rm q}}$, $b={v_{\rm q}}/{r_{\rm q}}$ are fixed so that the ansatz is consistent with the quenched-noise approximation. The emerging stationary variance $C_{-}(0)={r_{\rm q}}{g_{\rm q}}^{(\text{hand})}(0)-{r_{\rm q}}^{2}$ is, indeed, the stationary variance according to the quenched-noise approximation (in analogy with ${r_{\rm q}}$ in Eq. (40) in Sec. E):
${v_{\rm q}}({\mu_{\rm I}}):=$ ${r_{\rm w}}(\tilde{\mu})^{2}\right>_{{\cal N}(\tilde{\mu}|\mu_{I},{v_{\rm c}})}-{r_{\rm q}}^{2}.$ (70)
@@
${v_{\rm q}}({\mu_{\rm I}})+{r_{\rm q}}^{2}({\mu_{\rm I}})$ $={r_{\rm w}}({\mu_{\rm I}})^{2}\left[\frac{1}{1-2{v_{\rm c}}g^{\prime}_{1}({\mu_{\rm I}})}\right]^{1/2}\exp\left(\frac{2{v_{\rm c}}g_{1}^{2}({\mu_{\rm I}})}{1-2{v_{\rm c}}g^{\prime}_{1}({\mu_{\rm I}})}\right).$ (71)
@@
${\rm Var}\left[\frac{{n_{\rm R}}({T_{\rm c}})}{{T_{\rm c}}}\right]$ $\simeq\frac{{r_{\rm q}}}{{T_{\rm c}}}+\frac{2{\tau_{\rm c}}}{{T_{\rm c}}}\,{v_{\rm q}}.$ (72)
H.2 The variance of step-inhomogeneous Poisson processes
@@
where expectation values and variances refer to the annealed probability distributions of both the realisations of the Poisson spikes given $r_{\mu}$, and of the different instances of $r_{\mu}$. This is equivalent to suppose that one first samples $M$ independent $\mu\sim{\cal N}(\cdot;\mu_{I},v_{\rm c})$ and, for each of such $\mu$’s, one samples $n_{m}({T_{\rm s}})\sim{\rm Poisson}(\cdot;{T_{\rm s}}r_{\mu})$.
@@
Appendix I The encoding SNR
@@
Appendix J Dependence on ${\tau_{\rm c}}$ of the colored noise amplitude
We are interested in the dependence of the readout rate mean and variance, ${r_{\rm R}}^{(s)},{v_{\rm R}}^{(s)}$, on the encoding and decoding properties in the Valente model. In order to study the dependence on the correlation time ${\tau_{\rm c}}$, there are two possibilities (see references in [21]). The first one is to consider that the stationary noise amplitude ${v_{\rm c}}$ is constant with respect to ${\tau_{\rm c}}$ (being ${\sigma_{\rm c}}=(2{\tau_{\rm c}}{v_{\rm c}})^{1/2}$ consequently a function of ${\tau_{\rm c}}$). The second one is taking ${\sigma_{\rm c}}^{2}$ constant with respect to ${\tau_{\rm c}}$ (and, hence, ${\tau_{\rm c}}$-dependent stationary noise amplitude ${v_{\rm c}}={\sigma_{\rm c}}^{2}/(2{\tau_{\rm c}})$). The first scenario is suitable to address the ${\tau_{\rm c}}\to\infty$ limit of our problem since, if ${v_{\rm c}}$ is constant, the colored noise does have an influence for infinitely large ${\tau_{\rm c}}$ (indeed, $\text{Cov}\left(I(t),I(t^{\prime})\right)={v_{\rm c}}$, see Eq. (7)), while for ${\tau_{\rm c}}\to 0$ the coloured noise has no influence (indeed, $\text{Cov}\left(I(t),I(t^{\prime})\right)=0$). Oppositely, the constant-${\sigma_{\rm c}}$ scenario is the suitable one to consider the ${\tau_{\rm c}}\to 0$ regime of the encoding-readout model, since for ${\tau_{\rm c}}\to 0$ the colored noise term ${\sigma_{\rm c}}\xi(t)$ reduces to a delta-correlated noise ($\text{Cov}\left(I(t)I(t^{\prime})\right)=({\sigma_{\rm I}}^{2}+{\sigma_{\rm c}}^{2})\delta(t-t^{\prime})$, see Eq. (7)), while for infinite ${\tau_{\rm c}}$ it vanishes.
@@
We stress that, in any case, the constant ${v_{\rm c}}$ choice does not imply loss of generality, as it only determines the value of ${\sigma_{\rm c}}$ that we probe when varying ${\tau_{\rm c}}$. Equation (7) tells us that, the larger the value of ${\tau_{\rm c}}$, the larger must be ${\sigma_{\rm c}}$ in order for the colored noise term to have a significant impact in $I(t)$, hence in the readout neuron statistics. Taking constant ${v_{\rm c}}$, we automatically accounts for this issue. The constant-${\sigma_{\rm c}}$ scenario is left to future studies.
@@
Appendix K Underestimation of the rate variance
As already mentioned, we do not employ, in the main text, the finite-NIS correction algorithm to the variance, since it leads to large errors. Consequently, our estimations for ${v_{\rm R}}$ are slightly underestimated. This is, however, not a serious issue since, as we said, the overestimations in the numerator and denominator of the SNR in Fig. 8 partially cancel. In Fig. 20 we report the SNR but, differently from Fig. 8, we use the NIS-corrected estimation for the numerator, $\Delta{r_{\rm R}}$. In this situation, for our approximations to work well in the ${\tau_{\rm c}}\ll{\tau_{\rm m}}$ limit, one has to use the Poisson-Moreno approximation, but taking the numerical estimation for ${r_{\rm R}}$ (not the Moreno theoretical one) in the Poisson hypothesis: ${v_{\rm R}}=\tilde{r_{\rm R}}/{T_{\rm c}}$. This works, because the Poisson hypothesis $\tilde{r_{\rm R}}\simeq\tilde{v_{\rm R}}/{T_{\rm c}}$ still holds for finite ${n_{\rm is}}$, when both ${r_{\rm R}}$, ${v_{\rm R}}$ are underestimated. We consequently expect that the SNR in Fig. 20 is an overestimation of the SNR (since the numerator is NIS-corrected, while the denominator is underestimated), while the one in Fig. 8 is a more accurate estimation.
Figure 20: SNR vs ${\tau_{\rm c}}$, as in Fig. 8, but using the NIS-corrected estimator for the numerator, $\Delta{r_{\rm R}}$.
Appendix L SNR versus ${\mu_{\rm I}}^{(-)}$
See Figs. 21,22 for an analysis of how the SNR depends on ${\mu_{\rm I}}^{(\pm)}$ and on ${\sigma_{\rm c}}$. The figure illustrates how the behaviour in Fig. 9 (in which ${\mu_{\rm I}}^{(-)}$ is so low that ${r_{\rm R}}^{(-)}$ is negligible) changes for non-vanishing values of ${r_{\rm R}}^{(-)}$.
Figure 21: Readout rate mean, variance, and excess of SNR, as in Fig. 9, but for a higher ${\mu_{\rm I}}^{(-)}=340$ (signaled as a dashed vertical line). The rest of the model parameters are: ${\tau_{\rm c}}=0.1$, ${\tau_{\rm m}}=0.0025$, $w=0.0010$.
@@
Appendix M Theoretical SNR dependence on other variables
@@
In Figs. 23,24,25, the SNR has been estimated according to our analytical solutions: for low $\alpha_{\rm M}$, we use the Moreno et al solution for $\Delta{r_{\rm R}}$ (interpolated between the ${\tau_{\rm c}}\ll{\tau_{\rm m}}$ and ${\tau_{\rm c}}\gg{\tau_{\rm m}}$ solutions, as in [17]), and a Poisson-Moreno hypothesis for the denominator ${v_{\rm R}}^{(s)}\right>_{s}$ (i.e., supposing ${v_{\rm R}}={r_{\rm R}}/{T_{\rm c}}$, and the Moreno expression for ${r_{\rm R}}$); for large $\alpha_{\rm M}$, we use the quenched-noise approximation for both the numerator and the denominator of the SNR. In practice, whenever $\alpha_{\rm M}$ is larger than a threshold $\hat{\alpha}_{\rm M}=0.75$, we use the quenched-noise approximation. The green dashed line in Figs. 10,23,24,25 signals the $\alpha_{\rm M}=\hat{\alpha}_{\rm M}$ transitions in the $(\Delta{\mu_{\rm I}},{\tau_{\rm c}})$ plane. Although the Moreno and quenched-noise approximations are expected to hold only in the ${\tau_{\rm c}}\ll{\tau_{\rm m}}$ and ${\tau_{\rm c}}\gg{\tau_{\rm m}}$ limits, respectively, in practice they are relatively congruent around the $\alpha_{\rm M}=\hat{\alpha}_{\rm M}$ line for intermediate values of ${\tau_{\rm c}}$.
@@
Dependence on ${\alpha_{\rm V}}$. We report the dependence with ${\sigma_{\rm c}}$ (equivalently, ${\alpha_{\rm V}}$) in 25, for ${\alpha_{\rm V}}=0.25,0.9,1.5$. The largest probed value of ${\alpha_{\rm V}}$ is larger than one. While this is not, in principle, possible according to the definition of encoding noise correlations, in this section it is an effective way of increasing ${\sigma_{\rm c}}({\alpha_{\rm V}})=w{\sigma_{\rm V}}[N(1+(N-1){\alpha_{\rm V}}^{2})]^{1/2}$, above the value that it would have for ${\alpha_{\rm V}}=1$ (see Sec. C), and that could correspond to an equivalent increment in $N$, $w$ or ${\sigma_{\rm V}}$, keeping ${\alpha_{\rm V}}\leq 1$. First, for low enough ${\alpha_{\rm V}}$ (panel A), the amplitude of noise correlations is too low to imply a significant increment of SNR. For larger values of ${\alpha_{\rm V}}$, the non-monotonic dependence along both axes that we described in Sec. IV.2 emerge (panel B). For larger and larger ${\alpha_{\rm V}}$, and for a high enough value of ${\mu_{\rm I}}^{(-)}$, near enough $\hat{\mu_{\rm I}}^{(-)}$, the maximum of the SNR moves to the left, in the x-axis (panels B,C). This means that, at a fixed value of ${\mu_{\rm I}}^{(\pm)}$, increasing the value of ${\sigma_{\rm c}}$ may imply a lower $\hat{\mu_{\rm I}}^{(-)}$ and, consequently, a lower SNR, as we explained in Fig. 21.

### v2

@@ H.1 Sketch of the derivation of the temporal distribution function in the quenched-noise approximation
where $t,t^{\prime}:=\ell j,\ell j^{\prime}$, $P(\cdot)={\cal N}(\cdot;{\mu_{\rm I}},{v_{\rm c}})$ is the stationary distribution of input noises, and $P(\mu,t|\mu^{\prime},t^{\prime})$ is the Green’s function of an Ornstein-Uhlenbeck process: (mind $\xi$ in Eq. (5b)):
@@
An important remark on the validity limits of this derivation is that we expect Eq. 87 to be valid for ${\tau_{\rm m}}\ll t-t^{\prime}\ll{\tau_{\rm c}}$, i.e., to exhibit the validity limits of the general quenched-noise approximation. Afterwards, the analytical approximations to both nested integrations, that in ${r_{\rm w}}(\mu_{t})\right>_{P(\mu_{t}|\mu_{t^{\prime}})}$ (compacting the notation) and, afterwards, that in
@@
can be performed in the saddle-point approximation, for large ${v_{\rm c}}$, using, as in Appendix E.2, the variant of the Laplace integration method in Eq. (52). It is straightforward to arrive to an expression for ${g_{\rm q}}$ in such a quenched-noise approximation, for low enough ${v_{\rm c}}$. Interestingly, in the low- and large- $|t-t^{\prime}|$ limits, the ${r_{\rm R}}{g_{\rm q}}(t-t^{\prime})$ take, respectively, the values ${r_{\rm w}}^{2}(\mu)\right>_{P(\mu)}$ and ${r_{\rm w}}(\mu)\right>^{2}_{P(\mu)}$ (as expected, since for large $|t-t^{\prime}|$ we expect $\mu,\mu^{\prime}$ to be uncorrelated, and so ${r_{\rm w}}(\mu)$ and ${r_{\rm w}}(\mu^{\prime})$). In particular, for large enough $|t-t^{\prime}|$ (or, more precisely, neglecting terms of order $\omega^{2}$ in front of first-order terms in $\omega$), we get a simplified expression for ${g_{\rm q}}$:
@@
where $Q$ is a constant in time, depending on the white-LIF properties ${\mu_{\rm I}},{\sigma_{\rm I}},{\tau_{\rm m}}$, and on ${v_{\rm c}}$ (see the explicit expression in Eq. (99) below in Sec. H.2).
H.2 Evaluation of the nested integrals
@@
where ${t_{\rm w}}^{\prime},{t_{\rm w}}^{\prime\prime}$ are in Eq. (58), and where:
@@
Appendix I Hand-waving estimation for the rate variance in the quenched-noise approximation
I.1 Hand-waving ansatz for the temporal distribution function
We here present a hand-waving argument, consistent with the quenched-noise approximation, that leads to an alternative analytic expression for the rate variance of the LIF model with white+colored noise input. Interestingly, the hand-waving approximation for ${v_{\rm R}}$ and $g$ captures relatively well the behavior of the numeric estimation of these quantities, but worst than the full quenched-noise calculation in Sec. H, of which the hand-waving approximation is a further simplification (please, see Figs. 7,8).
@@ I.1 Hand-waving ansatz for the temporal distribution function
${g_{\rm q}}^{(\text{hand})}(t)=\frac{{u_{\rm q}}}{{r_{\rm q}}}e^{-|t|/{\tau_{\rm c}}}+{r_{\rm q}},$ (107)
where the coefficients $a={r_{\rm q}}$, $b={u_{\rm q}}/{r_{\rm q}}$ are fixed so that the ansatz is consistent with the quenched-noise approximation. The emerging stationary variance $C_{-}(0)={r_{\rm q}}{g_{\rm q}}^{(\text{hand})}(0)-{r_{\rm q}}^{2}$ is, indeed, the stationary variance according to the quenched-noise approximation (in analogy with ${r_{\rm q}}$ in Eq. (45) in Sec. E):
${u_{\rm q}}({\mu_{\rm I}}):=$ ${r_{\rm w}}(\tilde{\mu})^{2}\right>_{{\cal N}(\tilde{\mu}|\mu_{I},{v_{\rm c}})}-{r_{\rm q}}^{2}.$ (108)
@@
${u_{\rm q}}({\mu_{\rm I}})+{r_{\rm q}}^{2}({\mu_{\rm I}})$ $={r_{\rm w}}({\mu_{\rm I}})^{2}\left[\frac{1}{1-2{v_{\rm c}}g^{\prime}_{1}({\mu_{\rm I}})}\right]^{1/2}\exp\left(\frac{2{v_{\rm c}}g_{1}^{2}({\mu_{\rm I}})}{1-2{v_{\rm c}}g^{\prime}_{1}({\mu_{\rm I}})}\right).$ (109)
@@
${\rm Var}\left[\frac{{n_{\rm R}}({T_{\rm c}})}{{T_{\rm c}}}\right]$ $\simeq\frac{{r_{\rm q}}}{{T_{\rm c}}}+\frac{2{\tau_{\rm c}}}{{T_{\rm c}}}\,{u_{\rm q}}.$ (110)
I.2 The variance of step-inhomogeneous Poisson processes
@@
where expectation values and variances refer to the annealed probability distributions of both the realizations of the Poisson spikes given $r_{\mu}$, and of the different instances of $r_{\mu}$. This is equivalent to suppose that one first samples $M$ independent $\mu\sim{\cal N}(\cdot;\mu_{I},v_{\rm c})$ and, for each of such $\mu$’s, one samples $n_{m}({T_{\rm s}})\sim{\rm Poisson}(\cdot;{T_{\rm s}}r_{\mu})$.
@@
Appendix J Necessary conditions for the SNR enhancement
An upper bound in ${\mu_{\rm I}^{(+)}}$ for the SNR enhancement to occur. We will now argue that, since the SNR numerator is $\Delta{r_{\rm R}}\leq{r_{\rm R}}({\mu_{\rm I}}^{(+)})$, a necessary condition for the SNR to increase with the noise correlation strength is ${\mu_{\rm I}}^{(+)}<{\mu_{\rm I}}^{*}$. We will adopt the following double delta notation: $\delta$ for the increment when introducing the colored noise, ${\alpha_{\rm V}}>0$, while $\Delta$ for the $s=\pm$ change, with $\Delta{\mu_{\rm I}}={\mu_{\rm I}^{(+)}}-{\mu_{\rm I}^{(-)}}>0$. We focus in the regime when the white-noise LIF response curve ${r_{\rm w}}$ is convex, for low enough ${\mu_{\rm I}}<{\mu_{\rm I}}^{*}$, since it is the only one in which one observes a $\delta({\rm SNR})>0$. As we have argued above, in this regime it is $\delta{r_{\rm R}}>0$ (we can understand this fact thanks to the quenched-noise approximation, at least when ${\tau_{\rm c}}\gtrsim{\tau_{\rm m}}$). Moreover, we also know (again thanks to the quenched-noise hypothesis) that the increment in the SNR denominator in the presence of colored noise is positive $\delta\left<{v_{\rm R}}^{(s)}\right>_{s}$, since both terms of ${v_{\rm R}}$ in Eq. (11) increase with ${\alpha_{\rm V}}$. We conclude that, in the regime where one numerically observes $\delta({\rm SNR})>0$ ($\mu$ sufficiently lower than $\mu^{*}$), the condition $\delta({\rm SNR})>0$ implies $\delta\Delta{r_{\rm R}}>0$. Indeed, $\delta({\rm SNR})>0$ can be written as:
[…]
A lower upper bound in ${\mu_{\rm I}^{(+)}}$ for the SNR enhancement to occur. We can make the condition ${\mu_{\rm I}^{(+)}}\leq{\mu_{\rm I}}^{*}$ more stringent by imposing that, beyond the numerator, the SNR itself should increase. This lower upper bound for ${\mu_{\rm I}^{(+)}}$ is found by imposing the further condition that not only it has to be $\delta\Delta{r_{\rm R}}>0$, but $\delta({\rm SNR})>0$. First, we assume firing rates so low that, in the white-noise case for ${\alpha_{\rm V}}=0$, we can neglect the over-Poissonian deviations of ${r_{\rm R}}$ (mind that, for the white-noise model, this is verified in almost all cases probed in this article in which ${\alpha_{\rm V}}=0$, c.f. Fig. 8). Neglecting ${r_{\rm R}^{(-)}}$, and taking ${v_{\rm R}}{T_{\rm c}}\simeq{r_{\rm w}}$, the white-noise SNR is (we omit the ${T_{\rm c}}$ proportionality factors) ${\rm SNR}_{\rm w}\simeq(2{r_{\rm w}})^{1/2}$. Now, take ${\rm SNR}_{\rm q}\simeq{r_{\rm q}}^{(+)}({v_{\rm q}}^{(+)}/2)^{-1/2}$ (again neglecting ${r_{\rm R}^{(-)}}$), Taylor-expand our expression for ${r_{\rm q}}^{(+)}/{({v_{\rm q}}^{(+)})}^{1/2}$ (taking ${v_{\rm q}}$ from Eq. (109) in Appendix I), and stop at first order in ${v_{\rm c}}$. Equating $\text{SNR}_{\rm q}=\text{SNR}_{\rm w}$ results in the transcendental equation $g_{1}^{\prime}+g_{1}^{2}=2{\tau_{\rm c}}{r_{\rm w}}g_{1}^{2}$, whose solution is dubbed $\hat{\mu_{\rm I}}$. The numerator ${r_{\rm R}}^{(+)}$ and the denominator ${{v_{\rm R}}^{(+)}}^{1/2}$ both increase with ${\sigma_{\rm c}}$, so that both effects cancel at first order in ${v_{\rm c}}$. In fact, to order $o[{v_{\rm c}}]$, the condition $\text{SNR}_{\rm q}=\text{SNR}_{\rm w}$ translates into $g_{1}^{\prime}+g_{1}^{2}=2{\tau_{\rm c}}{r_{\rm w}}g_{1}^{2}$, and equation that does not depend on ${v_{\rm c}}$. The same happens to its solution $\hat{\mu_{\rm I}}$ (green line in Fig. 10).
We conclude that $\delta(\text{SNR})>0$ implies ${\mu_{\rm I}}^{(+)}<\hat{\mu_{\rm I}}$ for low enough ${v_{\rm c}}$ and ${\mu_{\rm I}^{(-)}}$ (since we have neglected ${r_{\rm R}^{(-)}}$). This condition can therefore be taken as a necessary condition on ${\mu_{\rm I}}^{(+)}$ for the SNR to increase in the presence of noise correlations, in the NERM, and for large enough ${\tau_{\rm c}}\gtrsim{\tau_{\rm m}}$, and low enough ${v_{\rm c}}$ and ${\mu_{\rm I}^{(-)}}$.
While we expect the upper bound ${\mu_{\rm I}}^{*}$ to hold always, the lower upper bound $\hat{\mu_{\rm I}}$ is therefore expected to hold for low enough values of ${\mu_{\rm I}^{(-)}}$ and ${v_{\rm c}}$. In Sec. IV.3 we have mentioned that not only ${\mu_{\rm I}}<{\mu_{\rm I}}^{*}$, but even the more stringent condition ${\mu_{\rm I}}<\hat{\mu_{\rm I}}$ is respected in all the figures in which we have evaluated the SNR. This is not very surprising, in the light of the fact that, for high enough values of ${\mu_{\rm I}^{(-)}}$ and of ${v_{\rm c}}$, also the condition $\delta(\text{SNR})>0$ tends to not be satisfied any more (as explained in Sec. IV.3). In practice, there are no cases, among those probed here, in which $\delta(\text{SNR})>0$ and, at the same time, the conditions low ${\mu_{\rm I}^{(-)}}$ and low ${v_{\rm c}}$ are not satisfied. Therefore, the upper bound $\hat{\mu_{\rm I}}$ tends to hold in all cases.
Appendix K The encoding SNR
@@
Appendix L Dependence on ${\tau_{\rm c}}$ of the colored noise amplitude
We are interested in the dependence of the readout rate mean and variance, ${r_{\rm R}}^{(s)},{v_{\rm R}}^{(s)}$, on the encoding and decoding properties in the Valente model. In order to study the dependence on the correlation time ${\tau_{\rm c}}$, there are two possibilities (see references in [21]). The first one is to consider that the stationary noise amplitude ${v_{\rm c}}$ is constant (where, throughout this section, by constant we mean independent of ${\tau_{\rm c}}$), being ${\sigma_{\rm c}}=(2{\tau_{\rm c}}{v_{\rm c}})^{1/2}$ consequently a function of ${\tau_{\rm c}}$. The second one is taking ${\sigma_{\rm c}}^{2}$ constant (and, hence, ${\tau_{\rm c}}$-dependent stationary noise amplitude ${v_{\rm c}}={\sigma_{\rm c}}^{2}/(2{\tau_{\rm c}})$). The first scenario is suitable to address the ${\tau_{\rm c}}\to\infty$ limit of our problem since, if ${v_{\rm c}}$ is constant, the colored noise does have an influence for infinitely large ${\tau_{\rm c}}$ (indeed, $\text{Cov}\left(I(t),I(t^{\prime})\right)={v_{\rm c}}$, see Eq. (7)), while for ${\tau_{\rm c}}\to 0$ the coloured noise has no influence (indeed, $\text{Cov}\left(I(t),I(t^{\prime})\right)=0$). Oppositely, the constant-${\sigma_{\rm c}}$ scenario is the suitable one to consider the ${\tau_{\rm c}}\to 0$ regime of the encoding-readout model, since for ${\tau_{\rm c}}\to 0$ the colored noise term ${\sigma_{\rm c}}\xi(t)$ reduces to a delta-correlated noise ($\text{Cov}\left(I(t)I(t^{\prime})\right)=({\sigma_{\rm I}}^{2}+{\sigma_{\rm c}}^{2})\delta(t-t^{\prime})$, see Eq. (7)), while for infinite ${\tau_{\rm c}}$ it vanishes (see Table 11).
@@
We stress that setting a constant value for ${v_{\rm c}}$ does not lead to a loss of generality, as it only determines the value of ${\sigma_{\rm c}}$ that we probe when varying ${\tau_{\rm c}}$. Equation (7) implies that, the larger the value of ${\tau_{\rm c}}$, the larger must be ${\sigma_{\rm c}}$ in order for the colored noise term to have a significant impact in $I(t)$, hence in the readout neuron statistics. Setting a constant value for ${v_{\rm c}}$ accounts for this issue. The constant-${\sigma_{\rm c}}$ scenario is left to future studies.
@@
Appendix M Underestimation of the rate variance
As already mentioned, we do not employ, in the main text, the finite-NIS correction algorithm to the variance, since it leads to large errors in the SNR. Consequently, our estimations for ${v_{\rm R}}$ are slightly underestimated. This is, however, not a serious issue since, as we said, the overestimations in the numerator and denominator of the SNR in Fig. 9 partially cancel. In Fig. 29 we report the SNR but, differently from Fig. 9, we use the NIS-corrected estimation for the numerator, $\Delta{r_{\rm R}}$. In this situation, for our approximations to work well in the ${\tau_{\rm c}}\ll{\tau_{\rm m}}$ limit, one has to use the Poisson-Moreno approximation, but taking the numerical estimation for ${r_{\rm R}}$ (not the Moreno theoretical one) in the Poisson hypothesis: ${v_{\rm R}}=\tilde{r_{\rm R}}/{T_{\rm c}}$. This works, because the Poisson hypothesis $\tilde{r_{\rm R}}\simeq\tilde{v_{\rm R}}/{T_{\rm c}}$ still holds for finite ${n_{\rm is}}$, when both ${r_{\rm R}}$, ${v_{\rm R}}$ are underestimated. We consequently expect that the SNR in Fig. 29 is an overestimation of the SNR (since the numerator is NIS-corrected, while the denominator is underestimated), while the one in Fig. 9 is a more accurate estimation.
Figure 29: SNR vs ${\tau_{\rm c}}$, as in Fig. 9, but using the NIS-corrected estimator for the numerator, $\Delta{r_{\rm R}}$. The thin dotted lines are the encoding signal-to-noise ratio in the unit time interval, Eq. (8). See the rest of model parameters in Table 4.
Appendix N SNR versus ${\mu_{\rm I}}^{(-)}$
See Figs. 30,31 for an analysis of how the SNR depends on ${\mu_{\rm I}}^{(\pm)}$ and on ${\sigma_{\rm c}}$. The figure illustrates how the behavior in Fig. 10 (in which ${\mu_{\rm I}}^{(-)}$ is so low that ${r_{\rm R}}^{(-)}$ is negligible) changes for non-vanishing values of ${r_{\rm R}}^{(-)}$.
Figure 30: Readout rate mean, variance, and excess of SNR, as in Fig. 10, but for a higher ${\mu_{\rm I}}^{(-)}=340$ (signaled as a dashed vertical line). The rest of the model parameters are: ${\tau_{\rm c}}=0.1$, ${\tau_{\rm m}}=0.0025$, $w=0.001$.
@@
Appendix O Theoretical SNR dependence on other variables
@@
In Figs. 32,33,34, the SNR has been estimated according to our analytical solutions: for low $\alpha_{\rm M}$, we use the Moreno et al solution for $\Delta{r_{\rm R}}$ (interpolated between the ${\tau_{\rm c}}\ll{\tau_{\rm m}}$ and ${\tau_{\rm c}}\gg{\tau_{\rm m}}$ solutions, as in [17]), and a Poisson-Moreno hypothesis for the denominator ${v_{\rm R}}^{(s)}\right>_{s}$ (i.e., supposing ${v_{\rm R}}={r_{\rm R}}/{T_{\rm c}}$, and the Moreno expression for ${r_{\rm R}}$); for large $\alpha_{\rm M}$, we use the quenched-noise approximation for both the numerator and the denominator of the SNR. In practice, whenever $\alpha_{\rm M}$ is larger than a threshold $\hat{\alpha}_{\rm M}=0.75$, we use the quenched-noise approximation. The green dashed line in Figs. 11,32,33,34 signals the $\alpha_{\rm M}=\hat{\alpha}_{\rm M}$ transitions in the $(\Delta{\mu_{\rm I}},{\tau_{\rm c}})$ plane. Although the Moreno and quenched-noise approximations are expected to hold only in the ${\tau_{\rm c}}\ll{\tau_{\rm m}}$ and ${\tau_{\rm c}}\gg{\tau_{\rm m}}$ limits, respectively, in practice they are relatively congruent around the $\alpha_{\rm M}=\hat{\alpha}_{\rm M}$ line for intermediate values of ${\tau_{\rm c}}$.
@@
Dependence on ${\alpha_{\rm V}}$. We report the dependence with ${\sigma_{\rm c}}$ (equivalently, ${\alpha_{\rm V}}$) in 34, for ${\alpha_{\rm V}}=0.25,0.9,1.5$. The largest probed value of ${\alpha_{\rm V}}$ is larger than one. While this is not, in principle, possible according to the definition of encoding noise correlations, in this section it is an effective way of increasing ${\sigma_{\rm c}}({\alpha_{\rm V}})=w{\sigma_{\rm V}}[N(1+(N-1){\alpha_{\rm V}}^{2})]^{1/2}$, above the value that it would have for ${\alpha_{\rm V}}=1$ (see Sec. C), and that could correspond to an equivalent increment in $N$, $w$ or ${\sigma_{\rm V}}$, keeping ${\alpha_{\rm V}}\leq 1$. First, for low enough ${\alpha_{\rm V}}$ (panel A), the amplitude of noise correlations is too low to imply a significant increment of SNR. For larger values of ${\alpha_{\rm V}}$, the non-monotonic dependence along both axes that we described in Sec. IV.3 emerge (panel B). For larger and larger ${\alpha_{\rm V}}$, and for a high enough value of ${\mu_{\rm I}}^{(-)}$, near enough $\hat{\mu_{\rm I}}^{(-)}$, the maximum of the SNR moves to the left, in the x-axis (panels B,C). This means that, at a fixed value of ${\mu_{\rm I}}^{(\pm)}$, increasing the value of ${\sigma_{\rm c}}$ may imply a lower $\hat{\mu_{\rm I}}^{(-)}$ and, consequently, a lower SNR, as we explained in Fig. 30.